OpenStudy (anonymous):

Solve by the linear combination method (with or without multiplication). 2x + 3y = 11 –4x + 2y = 2 HELP PLEASE GOT TEST NEEDS TO BE PASSED AND IM CRAP can i get the formla with the answer or jsut the answer :)

OpenStudy (dtan5457):

Can I solve this my method? I don't even know what linear combination is.

OpenStudy (dtan5457):

Well i'll leave you with this. I'd solve with elimination. your two equations are 2x+3y=11 -4x+2y=2 i'd divide the 2nd one by 2 to get -2x+y=1 2x+3y=11 -2x+y=1 subtract and you get 2y=10 y=5 plug back to any equation 2x+15=11 2x=-4 x=-2